how to find the value of 2 ^1/3.pls provide required steps

Grade:11

1 Answers

guddu gupta
3 years ago
2=e(ln2)( as a=e(lna))
we know that
ex= 1+ x + x2/2! + x3/3! ----(derived using taylor series)
2= e(ln2)x= 1 +(ln2)x + ((ln2)x)2/2! +((ln3)x)3/3! ----